OpenStudy (anonymous):

Write an equation for a circle with center at (2,3) and radius of 6

OpenStudy (anonymous):

(x-2)^2+(y-3)^2=36

OpenStudy (anonymous):

Equation of circle is given by: \[(x-h)^2 + (y-k)^2 = r^2\] Where, (h,k) = center and r = radius..

OpenStudy (anonymous):

thnks u guys

OpenStudy (anonymous):

So according to this: \[(x-2)^2 + (y-3)^2 = 6^2\] \[(x-2)^2 + (y-3)^2 = 36\]

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

thnk u i had to use an example so i wud kno how to do the rest so thnks

OpenStudy (anonymous):

Write an equation for a circle with center at (-2,7) and diameter of 14. i got (x-2)^2+(y-7)^2=196 but they said it was wrong @waterineyes

OpenStudy (anonymous):

@waterineyes ?

OpenStudy (anonymous):

You are given with diameter.. firstly find the radius from it: 2*r = diameter 2*r = 14 r = 7 S0, on RHS 49 will come..

OpenStudy (anonymous):

im lost so it ='s 49 not 196?

OpenStudy (anonymous):

Yes you have to calculate radius first..

OpenStudy (anonymous):

but i thought tht 14 wud be ^2?

OpenStudy (anonymous):

No that is the diameter.. You have to half it when diameter is given..

OpenStudy (anonymous):

but u didnt half 6

OpenStudy (anonymous):

on the other problem

OpenStudy (anonymous):

ohhhhhhh nevermind 6 was the radius

OpenStudy (anonymous):

nd 14 is the diameter

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

lol im sorry i get it

OpenStudy (anonymous):

it still says im wrong even with the 49???